OpenStudy (anonymous):

how do I get the derivative of y=(1-sinx)^(-1/2)

OpenStudy (anonymous):

use the chain rule

OpenStudy (anonymous):

the derivative of \(\sqrt{x}\) is \(\frac{1}{2\sqrt{x}}\)

OpenStudy (anonymous):

the derivative of sin(x) is cos(x)

OpenStudy (anonymous):

the derivative of \(x^a\) is \(ax^{a-1}\)

OpenStudy (anonymous):

use all these concepts and apply the chain rule

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

Np ;)